OpenStudy (anonymous):

Which number sentence matches the question below? For how many days did Curtis practice juggling, if he practiced for 200 hours in all and for 10 hours each day? 10 ÷ n = 200 200 ÷ n = 10 n ÷ 200 = 10 n ÷ 10 = 200

OpenStudy (anonymous):

200 ÷ n = 10

OpenStudy (anonymous):

because 200 ÷ 10 = 20 so... 200 ÷ 20 = 10 so n = 20 s0 200 ÷ n = 10